OpenStudy (anonymous):

Choose the correct simplification of the expression (d3)5.

OpenStudy (anonymous):

i got d15 but am not sure if thats correct

OpenStudy (phi):

yes, that is correct. btw, you should write it as d^15 (the uparrow means the 15 is an exponent.) \[ (d^3)^5 \text{ means } (d^3)(d^3)(d^3)(d^3)(d^3) \] d^3 times itself 5 times but d^3 means d*d*d so we get \[ (d^3)^5 = (d\cdot d\cdot d) (d\cdot d\cdot d) (d\cdot d\cdot d) (d\cdot d\cdot d) (d\cdot d\cdot d)\] or \[ (d^3)^5 = d\cdot d\cdot d\cdot d\cdot d\cdot d\cdot d\cdot d\cdot d\cdot d\cdot d\cdot d\cdot d\cdot d\cdot d\] now we can count up the d's. of course it is faster to just multiply 3*5=15 (a good trick to remember)

OpenStudy (anonymous):

thanks, that really helpful ........... ill save this in my notes